• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Calarco

    Log entry time 15:57:39 on October28,2001

    Entry number 75321

    keyword=Shift summary

    Altogether, not a productive shift. It started off with cryogenic problems
    in the HRSs due to a communications failure caused by IOC trips. The
    solution was to take all magnets to 0, reset everything, and bring them
    up again (see entries by Simon, no. 75303, and myself, no. 75309).
    After that we finished the runs at the A+ kinematics between 11:30 and
    12:00. We then proceeded to change HRS-right to 74.82 deg and take
    collimators to open position. Since then, the beam has been sparse at
    best. We started runs for the G+ kinematics at 13:30. The following
    runs have been completed:
    .
    Run 1331 with 0.8M events in 34 uA-hrs, terminated by Q2 trip
    Run 1332 with 1.0M events in 44 uA-hrs, terminated normally
    .
    The number of uA-hrs is calculated by I*N/(r*3600) where I is the
    beam current when on, typically 76-80 uA, N is the number of events,
    typically 1M for a good run, r is the CODA event rate, typically 500 Hz
    for the G+ runs.